# Collective dynamics versus entanglement in quantum battery performance

### v1

@@ I Introduction
Quantum batteries (QBs) are devices engineered to store and deliver energy through controlled quantum dynamics and have emerged as a central platform in quantum thermodynamics and quantum technologies. Importantly, QBs can be realized using quantum many-body systems, such as interacting spin chains, where collective effects, many-body interactions, and correlations can enhance charging power and lead to nontrivial scaling behaviors [alicki2013entanglement, binder2015quantacell, campaioli2017enhancing, ferraro2018high, friis2018precision, andolina2019quantum, andolina2019extractable, campaioli2018quantum, le2018spin, lyon2020many, kurizki2020quantum, rodriguez2021collective, zhang2022longrange, giuliani2022collective, zanardi2022scrambling]. Beyond conventional spin-chain architectures, QBs have been investigated in topological systems as well as in various spin-based models subject to different types of external driving or interactions, including Floquet-engineered setups, kicked-spin models, and long-range interacting frameworks. These studies highlight the roles of periodic driving, operator spreading, and topological protection in the storage and transfer of energy [lu2025topological, Puri2024Floquet, Romero2025].
@@
Moreover, QBs have been extensively investigated in open quantum system scenarios, where environmental effects such as dissipation, decoherence, and engineered reservoirs strongly influence charging efficiency, stability, and extractable work. These studies reveal phenomena including noise-assisted and dissipative charging, reservoir-mediated energy transfer, non-Markovian memory effects, thermal and composite environment influences, and fundamental power bounds [Ferraro2018, Andolina2019, GarciaPintos2020, Sharma2022, Quach2020, Tabesh2020EnvMediated, Morrone2022NonMarkov, EnhanceOpenQB2023, Yao2025Reservoir, QuantumErgotropy2025, NonMarkovNSpin2025, Xu2023CompositeEnv, Zhang2024OpenInteractive, Zhang2024OpenEnvPerformance, EntropicUncertaintyQB2022]. Collectively, these works emphasize that both the internal structure of the quantum battery and the surrounding environment critically determine charging performance.
@@
From a structural perspective, quantum battery architectures are commonly categorized into two paradigms: (i) interacting batteries charged by noninteracting systems [le2018spin, ghosh2020enhancement, zhao2021quantum], and (ii) noninteracting batteries charged by interacting many-body chargers [rossini2020quantum, ferraro2018high, julia2020bounds]. In both cases, charging performance depends sensitively on the Hamiltonian structure, interaction range, and normalization, with rigorous bounds linking charging power to the Hamiltonian norm [julia2020bounds, abdelkhalek2019].
@@
A central goal in quantum battery research is to identify genuinely quantum mechanisms such as coherence, entanglement, and collective dynamics that can enhance stored energy, extractable work, and especially charging power beyond classical limits [baumgratz2014quantifying, horodecki2009quantum]. Early studies demonstrated that global collective driving can enhance instantaneous charging power relative to parallel local protocols [alicki2013entanglement, binder2015quantacell, campaioli2017enhancing], motivating the idea that entanglement and multipartite correlations might serve as key resources for fast charging. However, subsequent investigations clarified that entanglement is neither necessary nor sufficient for charging enhancement [andolina2019quantum, andolina2019extractable, julia2020bounds, kamin2020entanglement, shi2022entanglement], as maximal charging power is often reached prior to the development of strong multipartite correlations [kamin2020entanglement, shi2022entanglement, perarnau2015extractable].
Recent attention has focused on the role of dynamical complexity, including quantum chaos, operator spreading, and scrambling. Studies based on Floquet engineering, kicked-Ising models, and long-range interactions indicate that rapid operator growth can significantly boost charging performance [Puri2024Floquet, Romero2025]. Likewise, works on information scrambling show that fast propagation of quantum information, characteristic of nonintegrable systems, can optimize energy transfer in many-body setups [nahum2018operator, vonkeyserlingk2018operator, liu2018scrambling, swingle2018unscrambling], suggesting that dynamical properties, rather than static entanglement alone, are decisive in determining charging power.
@@
To study correlations among three subsystems simultaneously, we use the tripartite mutual information (TMI), defined as [iyoda2018scrambling, seshadri2018tripartite, aleiner2016microscopic, xu2022operator, zhang2020information, hosur2016chaos, von2018scrambling, Shukla2025ScramblingIsing]
@@
While BEE measures correlations for a single partition, the average bipartite entanglement entropy (ABEE) provides a global view of entanglement across the system [Polkovnikov2011, Riera2012]. It is computed by averaging the BEE over all contiguous bipartitions:

### v2

@@ I Introduction
Quantum batteries (QBs) are devices engineered to store and deliver energy through controlled quantum dynamics and have emerged as a central platform in quantum thermodynamics and quantum technologies. Importantly, QBs can be realized using quantum many-body systems, such as interacting spin chains, where collective effects, many-body interactions, and correlations can enhance charging power and lead to nontrivial scaling behaviors [alicki2013entanglement, binder2015quantacell, campaioli2017enhancing, ferraro2018high, andolina2019quantum, andolina2019extractable, le2018spin, rodriguez2021collective]. Beyond conventional spin-chain architectures, QBs have been investigated in topological systems as well as in various spin-based models subject to different types of external driving or interactions, including Floquet-engineered setups, kicked-spin models, and long-range interacting frameworks. These studies highlight the roles of periodic driving, operator spreading, and topological protection in the storage and transfer of energy [Puri2024FloquetManyBody, lu2025topological, Puri2024FloquetManyBody, Romero2025, sahoo2025power].
@@
Moreover, QBs have been extensively investigated in open quantum system scenarios, where environmental effects such as dissipation, decoherence, and engineered reservoirs strongly influence charging efficiency, stability, and extractable work. These studies reveal phenomena including noise-assisted and dissipative charging, reservoir-mediated energy transfer, non-Markovian memory effects, thermal and composite environment influences, and fundamental power bounds [Ferraro2018, Andolina2019, GarciaPintos2020, Quach2020, Tabesh2020EnvMediated, Morrone2022NonMarkov, EnhanceOpenQB2023, Yao2025Reservoir, QuantumErgotropy2025, NonMarkovNSpin2025, Xu2023CompositeEnv, Zhang2024OpenInteractive, Zhang2024OpenEnvPerformance, EntropicUncertaintyQB2022, liu2025open, mathew2025super]. Collectively, these works emphasize that both the internal structure of the quantum battery and the surrounding environment critically determine charging performance.
@@
From a structural perspective, quantum battery architectures are commonly categorized into two paradigms: (i) interacting batteries charged by noninteracting systems [le2018spin, ghosh2020enhancement, zhao2021quantum], and (ii) noninteracting batteries charged by interacting many-body chargers [rossini2020quantum, ferraro2018high, julia2020bounds]. In both cases, charging performance depends sensitively on the Hamiltonian structure, interaction range, and normalization, with rigorous bounds linking charging power to the Hamiltonian norm [julia2020bounds].
@@
A central goal in quantum battery research is to identify genuinely quantum mechanisms such as coherence, entanglement, and collective dynamics that can enhance stored energy, extractable work, and especially charging power beyond classical limits [gyhm2024beneficial]. Early studies demonstrated that global collective driving can enhance instantaneous charging power relative to parallel local protocols [alicki2013entanglement, binder2015quantacell, campaioli2017enhancing], motivating the idea that entanglement and multipartite correlations might serve as key resources for fast charging. However, subsequent investigations clarified that entanglement is neither necessary nor sufficient for charging enhancement [andolina2019quantum, andolina2019extractable, julia2020bounds, kamin2020entanglement, shi2022entanglement], as maximal charging power is often reached prior to the development of strong multipartite correlations [kamin2020entanglement, shi2022entanglement, perarnau2015extractable].
@@
To study correlations among three subsystems simultaneously, we use the tripartite mutual information (TMI), defined as [iyoda2018scrambling, seshadri2018tripartite, Shukla2025ScramblingIsing]
@@
While BEE measures correlations for a single partition, the average bipartite entanglement entropy (ABEE) provides a global view of entanglement across the system. It is computed by averaging the BEE over all contiguous bipartitions: